# Post-detection inference for sequential changepoint localization

### v3

@@ B.2 Another concrete example: Poisson rate change problem
$\mathcal{S}^{\prime}_{n}=\{\theta:n\theta-\sum_{i=t}^{t+n-1}X_{i}\log\theta\leq\log\left(\frac{1}{\beta}\right)+\log I(c,c\theta)-\log I(n+c,\sum_{i=t}^{t+n-1}X_{i}+c\theta)\}\cap\Theta_{1},$ (B.13)
[…]
@@
$Y_{n}^{t}(X_{n}^{j},\theta):=\begin{cases}\sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta_{0}/\lambda_{t}),\text{ if }n<t\\ \sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta/\lambda_{t}),\text{ otherwise. }\end{cases}$ (B.14)
@@
We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau})$, as defined in Section 3, with $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$. Let $\tau^{\theta}_{j}$ be the detection time for the sequence $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$ and $M_{t,\infty}^{j}(\theta)=M_{t}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{\tau^{\theta}_{j}}^{t}(X_{\tau^{\theta}_{j}}^{j},\theta))$. Suppose, we can find $t_{1}^{j},t_{2}^{j}\in\mathbb{N}$, such that $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ (these are easy to compute for specific detection algorithms, such as CUSUM or SR type detectors, as we explain in Remark B.3) and they are finite. So, we can choose $L=\infty$.
@@ Appendix C Computationally and statistically efficient subroutines in Algorithm 3
$L_{t}(\theta_{0},\theta_{1};X_{1},\cdots,X_{n})=\prod_{i=1}^{t-1}f_{\theta_{0}}(X_{i})\prod_{i=t}^{n}f_{\theta_{1}}(X_{i}).$ (C.1)
@@
$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\hat{\theta}_{0,1:j-1}}(X_{i})}.$ (C.2)
@@
$M_{t}^{(n)}(X_{1},\cdots,X_{n}):=\displaystyle\max_{1\leq i\leq n}\frac{L_{i}(\hat{\theta}_{0,1:i-1},\hat{\theta}_{1,i:n};X_{1},\cdots,X_{n})}{L_{t}(\hat{\theta}_{0,1:t-1},\hat{\theta}_{1,t:n};X_{1},\cdots,X_{n})}.$ (C.3)
@@
$M_{t}=\begin{cases}\frac{L_{\hat{T}}(\hat{\theta}_{0,1:\hat{T}-1},\hat{\theta}_{1,\hat{T}:\tau};X_{1},\cdots,X_{\tau})}{L_{t}(\hat{\theta}_{0,1:t-1},\hat{\theta}_{1,t:\tau};X_{1},\cdots,X_{\tau})},&\text{if }t\leq\tau<\infty,\\ -\infty,&\text{if }t>\tau\text{ or }\tau=\infty.\\ \end{cases}$ (C.4)
As in Appendix B, we assume that we have a fixed element $\theta^{\prime\prime}$ in the parameter space and a function $G$, such that $X\sim F_{\theta^{\prime\prime}}$ implies $G(X,\theta)\sim F_{\theta},\forall\theta\in\Theta_{0}\cup\Theta_{1}$. Therefore, for each $j=1,\cdots,B$, we generate data sequence $X_{1}^{j},X_{2}^{j}\cdots\stackrel{{\scriptstyle iid}}{{\sim}}F_{\theta^{\prime\prime}}$ only once. Then, for each $t\in\{1,\cdots,\tau\}$ and $(\theta\times\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$, we obtain sequences $\{Y_{n}^{t}(X_{n}^{j},\theta,\theta^{\prime})\}_{n}$ having joint distribution $\mathbb{P}_{\theta,t,\theta^{\prime}},$ by applying the function $G(.,\theta)$ to $X_{1}^{j},\cdots,X_{t-1}^{j}$ and $G(.,\theta^{\prime})$ to $X_{t}^{j},X_{t+1}^{j},\cdots$, i.e.,
$Y_{n}^{t}(x,\theta,\theta^{\prime}):=\begin{cases}G(x,\theta),&\text{if }n<t\\ G(x,\theta^{\prime}),&\text{ otherwise. }\end{cases}$ (C.5)
@@
$U_{t,L}^{j}:=\begin{cases}-\infty,\leavevmode\nobreak\ \leavevmode\nobreak\ \quad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{ if }t>t_{2}^{j}\\ \displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),\\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{if }t\leq t_{2}^{j}\leq L\\ \infty,\leavevmode\nobreak\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{ if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L.\end{cases}$ (C.6)
@@
$U_{t,\infty}^{j}:=\begin{cases}-\infty,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{if }t>t_{2}^{j}\text{ or }t_{1}^{j}=\infty\\ \displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),\\ \leavevmode\nobreak\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{otherwise.}\end{cases}$ (C.7)
@@
$\mathcal{C}^{\prime\prime}_{1}=\{t\in\{2,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},U_{t,L}^{1},\cdots,U_{t,L}^{B})\}\text{ and}$ (C.8)
@@
Corollary C.1.
$\mathcal{C}^{\prime\prime}_{1}$ defined in (C.8) satisfies the same coverage guarantees as $\mathcal{C}_{1}$ in Theorem 4.5.
@@
C.1 A concrete example: Gaussian mean shift problem
@@
$\mathcal{S}_{t-1}=\left[\bar{X}_{1:t-1}-\frac{q_{\gamma/2}}{\sqrt{t-1}},\bar{X}_{1:t-1}+\frac{q_{\gamma/2}}{\sqrt{{t-1}}}\right]\cap\Theta_{0},$ (C.9)
and the $1-\beta$ confidence sequence for $\theta_{1}$, as defined in (B.9).
@@
$Y_{n}^{j}(\theta,\theta^{\prime})=\begin{cases}\theta+\epsilon_{n}^{j},\text{ if }n<t\\ \theta^{\prime}+\epsilon_{n}^{j},\text{ otherwise. }\end{cases}$ (C.10)
We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau};\tau)$, as defined in (C.4), with $\hat{\theta}_{0,1:t-1}=\frac{1}{t-1}\sum_{k=1}^{t-1}X_{k}$ and $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$. Note that, $t_{1}^{j},t_{2}^{j}$ are easy to compute exactly even with $L=\infty$ for some specific detection algorithms; see Remark C.2. For $t_{2}^{j}<t,U_{t,\infty}^{j}=0$ and for $t_{2}^{j}\geq t,U_{t,\infty}^{j}$ can be simplified for our Gaussian setting as below:
@@
$\begin{cases}\displaystyle\sum_{k=i}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}-\frac{t^{\prime}-t+1}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}\\ \quad+\sum_{k=t}^{t^{\prime}}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}+\frac{t-i}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}+\text{constant},&\text{if }i<t,\\[8.0pt] \text{constant},&\text{if }i=t,\\[8.0pt] \displaystyle\sum_{k=1}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}-\frac{i-t}{i-1}(\theta_{1}-\theta_{0})\right)^{2}\\ \quad+\sum_{k=t}^{i-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}+\frac{t-1}{i-1}(\theta_{1}-\theta_{0})\right)^{2}+\text{constant},&\text{if }i>t,\end{cases}$ (C.11)
over $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$. Here, constant means that it does not involve $\theta,\theta^{\prime}$. Note that $\delta_{t^{\prime}}^{i}:=\theta_{1,t^{\prime}}^{i}-\theta_{0,t^{\prime}}^{i}$ have simple closed-form expressions since (C.1) is a quadratic function of $\theta_{1}-\theta_{0}$. Then, we can take $(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i})$ to be any pair $(\theta,\delta_{t^{\prime}}^{i}+\theta)\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$.
Remark C.2.
[…]
$\tau_{\text{wcs-ripr}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (C.12)
$\tau_{\text{wsr-ripr}}=\inf\left\{n\in\mathbb{N}:\sum_{j=1}^{n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (C.13)
@@
The Gaussian example we provided serves as a concrete and illustrative case to clarify the methodology. Note that $\mathcal{C}^{\prime\prime}$ in (C.8) is computable if similar conditions as listed in Section B.1 are met. For instance, one can check that these conditions hold for analogous Laplace and exponential scale distributions. Another concrete example with the Poisson distribution is in Section G.3.
C.2 Another concrete example: Poisson rate change problem

### v4

@@ C.2 Another concrete example: Poisson rate change problem
$\mathcal{S}^{\prime}_{n}=\{\theta:n\theta-\sum_{i=t}^{t+n-1}X_{i}\log\theta\leq\log\left(\frac{1}{\beta}\right)+\log I(c,c\theta)-\log I(n+c,\sum_{i=t}^{t+n-1}X_{i}+c\theta)\}\cap\Theta_{1},$ (S13)
where $c$ is any constant and $I(a,b)=\int_{-\infty}^{\infty}e^{-ae^{x}+bx}dx$.
@@
$Y_{n}^{t}(X_{n}^{j},\theta):=\begin{cases}\sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta_{0}/\lambda_{t}),\text{ if }n<t\\ \sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta/\lambda_{t}),\text{ otherwise. }\end{cases}$ (S14)
@@
We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau})$, as defined in Section 3, with $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$. Let $\tau^{\theta}_{j}$ be the detection time for the sequence $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$ and $M_{t,\infty}^{j}(\theta)=M_{t}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{\tau^{\theta}_{j}}^{t}(X_{\tau^{\theta}_{j}}^{j},\theta))$. Suppose, we can find $t_{1}^{j},t_{2}^{j}\in\mathbb{N}$, such that $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ (these are easy to compute for specific detection algorithms, such as CUSUM or SR type detectors, as we explain in Remark C.3) and they are finite. So, we can choose $L=\infty$.
@@
Appendix D Computationally and statistically efficient subroutines in Algorithm 3
@@
$L_{t}(\theta_{0},\theta_{1};X_{1},\cdots,X_{n})=\prod_{i=1}^{t-1}f_{\theta_{0}}(X_{i})\prod_{i=t}^{n}f_{\theta_{1}}(X_{i}).$ (S1)
@@
$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\hat{\theta}_{0,1:j-1}}(X_{i})}.$ (S2)
@@
$M_{t}^{(n)}(X_{1},\cdots,X_{n}):=\displaystyle\max_{1\leq i\leq n}\frac{L_{i}(\hat{\theta}_{0,1:i-1},\hat{\theta}_{1,i:n};X_{1},\cdots,X_{n})}{L_{t}(\hat{\theta}_{0,1:t-1},\hat{\theta}_{1,t:n};X_{1},\cdots,X_{n})}.$ (S3)
@@
[…]
As in Appendix C, we assume that we have a fixed element $\theta^{\prime\prime}$ in the parameter space and a function $G$, such that $X\sim F_{\theta^{\prime\prime}}$ implies $G(X,\theta)\sim F_{\theta},\forall\theta\in\Theta_{0}\cup\Theta_{1}$. Therefore, for each $j=1,\cdots,B$, we generate data sequence $X_{1}^{j},X_{2}^{j}\cdots\stackrel{{\scriptstyle iid}}{{\sim}}F_{\theta^{\prime\prime}}$ only once. Then, for each $t\in\{1,\cdots,\tau\}$ and $(\theta\times\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$, we obtain sequences $\{Y_{n}^{t}(X_{n}^{j},\theta,\theta^{\prime})\}_{n}$ having joint distribution $\mathbb{P}_{\theta,t,\theta^{\prime}},$ by applying the function $G(.,\theta)$ to $X_{1}^{j},\cdots,X_{t-1}^{j}$ and $G(.,\theta^{\prime})$ to $X_{t}^{j},X_{t+1}^{j},\cdots$, i.e.,
$Y_{n}^{t}(x,\theta,\theta^{\prime}):=\begin{cases}G(x,\theta),&\text{if }n<t\\ G(x,\theta^{\prime}),&\text{ otherwise. }\end{cases}$ (S5)
@@
$U_{t,L}^{j}:=\begin{cases}-\infty,\penalty 10000\ \penalty 10000\ \quad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{ if }t>t_{2}^{j}\\ \displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),\\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{if }t\leq t_{2}^{j}\leq L\\ \infty,\penalty 10000\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{ if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L.\end{cases}$ (S6)
@@
$U_{t,\infty}^{j}:=\begin{cases}-\infty,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{if }t>t_{2}^{j}\text{ or }t_{1}^{j}=\infty\\ \displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),\\ \penalty 10000\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{otherwise.}\end{cases}$ (S7)
@@
$\mathcal{C}^{\prime\prime}_{1}=\{t\in\{2,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},U_{t,L}^{1},\cdots,U_{t,L}^{B})\}\text{ and}$ (S8)
@@
Corollary D.1.
$\mathcal{C}^{\prime\prime}_{1}$ defined in (S8) satisfies the same coverage guarantees as $\mathcal{C}_{1}$ in Theorem 4.4.
@@
D.1 A concrete example: Gaussian mean shift problem
@@
$\mathcal{S}_{t-1}=\left[\bar{X}_{1:t-1}-\frac{q_{\gamma/2}}{\sqrt{t-1}},\bar{X}_{1:t-1}+\frac{q_{\gamma/2}}{\sqrt{{t-1}}}\right]\cap\Theta_{0},$ (S9)
and the $1-\beta$ confidence sequence for $\theta_{1}$, as defined in (S9).
@@
$Y_{n}^{j}(\theta,\theta^{\prime})=\begin{cases}\theta+\epsilon_{n}^{j},\text{ if }n<t\\ \theta^{\prime}+\epsilon_{n}^{j},\text{ otherwise. }\end{cases}$ (S10)
We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau};\tau)$, as defined in (S4), with $\hat{\theta}_{0,1:t-1}=\frac{1}{t-1}\sum_{k=1}^{t-1}X_{k}$ and $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$. Note that, $t_{1}^{j},t_{2}^{j}$ are easy to compute exactly even with $L=\infty$ for some specific detection algorithms; see Remark D.2. For $t_{2}^{j}<t,U_{t,\infty}^{j}=0$ and for $t_{2}^{j}\geq t,U_{t,\infty}^{j}$ can be simplified for our Gaussian setting as below:
@@
$\begin{cases}\displaystyle\sum_{k=i}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}-\frac{t^{\prime}-t+1}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}\\ \quad+\sum_{k=t}^{t^{\prime}}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}+\frac{t-i}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}+\text{constant},&\text{if }i<t,\\[8.0pt] \text{constant},&\text{if }i=t,\\[8.0pt] \displaystyle\sum_{k=1}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}-\frac{i-t}{i-1}(\theta_{1}-\theta_{0})\right)^{2}\\ \quad+\sum_{k=t}^{i-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}+\frac{t-1}{i-1}(\theta_{1}-\theta_{0})\right)^{2}+\text{constant},&\text{if }i>t,\end{cases}$ (S11)
over $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$. Here, constant means that it does not involve $\theta,\theta^{\prime}$. Note that $\delta_{t^{\prime}}^{i}:=\theta_{1,t^{\prime}}^{i}-\theta_{0,t^{\prime}}^{i}$ have simple closed-form expressions since (D.1) is a quadratic function of $\theta_{1}-\theta_{0}$. Then, we can take $(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i})$ to be any pair $(\theta,\delta_{t^{\prime}}^{i}+\theta)\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$.
Remark D.2.
For weighted CUSUM or SR detectors, finding $t_{1}^{j},t_{2}^{j}$ is often straightforward. So, for some weight $W(\theta)$ such that $\int_{\theta\in\Theta_{1}}dW(\theta)=1$, the weighted CUSUM or SR detectors (with the Reverse Information Projection (RIPr) [17, 8, 16] being the representative of the pre-change class) are defined as
$\tau_{\text{wcs-ripr}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (S12)
$\tau_{\text{wsr-ripr}}=\inf\left\{n\in\mathbb{N}:\sum_{j=1}^{n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (S13)
@@
The Gaussian example we provided serves as a concrete and illustrative case to clarify the methodology. Note that $\mathcal{C}^{\prime\prime}$ in (S8) is computable if similar conditions as listed in Section C.1 are met. For instance, one can check that these conditions hold for analogous Laplace and exponential scale distributions. Another concrete example with the Poisson distribution is in Section H.4.
D.2 Another concrete example: Poisson rate change problem